# Event-Based Simulation of Stochastic Memristive Devices for Neuromorphic Computing

### v1

@@ Abstract
In this paper, we build a general model of memristors suitable for the simulation of event-based systems, such as hardware spiking neural networks, and more generally, neuromorphic computing systems. We extend an existing general model of memristors - the Generalised Metastable Switch Model [1] - to an event-driven setting, eliminating errors associated discrete time approximation, as well as offering potential improvements in terms of computational efficiency for simulation. We introduce the notion of a volatility state variable, to allow for the modelling of memory-dependent and dynamic switching behaviour, succinctly capturing and unifying a variety of volatile phenomena present in memristive devices, including state relaxation, structural disruption, Joule heating, and drift acceleration phenomena. We supply a drift dataset for titanium dioxide memristors and introduce a linear conductance model to simulate the drift characteristics, motivated by a proposed physical model of filament growth. We then demonstrate an approach for fitting the parameters of the event-based model to the drift model.
@@ I Introduction
The modelling of memristive devices is an important component of the development of novel architectures and systems based on these devices, including neuromorphic and event-based computing systems. Models can be classified along several dimensions: their generality, whether they are data-driven or physics based (or both), the inclusion of stochastic or non-ideal behaviour, and suitability for simulation in terms of their computational efficiency (compactness) in the relevant simulation context. The last of these dimensions can be thought of as key - since the goal of a model is to enable the prediction of behaviour so as to allow for the design of strategies to utilise the behaviour of the devices to our advantage. The specific goal of our modelling approach is the development of event-based neuromorphic computing simulations. Accordingly, this paper aims to extend existing modelling approaches for memristive devices to an event-based setting to allow the models to be used for neuromorphic computing simulations, by introducing general modelling principles, applicable to a broad range of memristive devices.
Figure 1: A model’s use is governed by its predictive ability in a particular setting. We focus on memristor modelling appropriate for use in a neuromorphic computing setting, introducing event-based simulation and the use of a volatility state variable to enable the model to be used to simulate key behaviours in this setting.
I-A Memristors
The term memristor has come to encompass a great many types of device, since its proposition as the fourth fundamental passive circuit element by Chua [2]. The original formulation proposed an equation linking the charge to the flux through a (polynomial) relationship, which resulted in an instantaneous resistance determined by the ratio of the voltage to the current, and dependent (in some way) on the instantaneous charge present in the device. The rate of change of the device’s resistive state was purely a function of the state itself and the magnitude of the voltage applied across the device (or the current flowing into it).
To expand the scope of memristors, Chua & Kang published a paper introducing the notion of memristive systems [3]. Memristive systems were defined purely mathematically, with no restrictions on the nature of the physical quantities, but only the relationship between them, also allowing for one or more state variables to exist in the device. Despite its generality, this model was unable to capture autonomous state evolution dynamics (later addressed by other models [4]), and was far removed from the physical world, by design, to allow it to be as general as possible, limiting its use for modelling real devices.
A change in the definition of the term “memristor” was later ignited by the identification of reasonably reliable resistive switching behaviour in a titanium dioxide thin film device [5], and subsequently fuelled by the broadening of the definition by Chua [6]. The types of device that have been described as memristors, commonly grouped according to the switching mechanism, include Phase Change Memory (PCM), Electrochemical Metallization (ECM), Valence Change Mechanism (VCM), Thermochemical Mechanism (TCM), Electronic (including Mott insulator and electron trapping devices), magnetoresistive memories (i.e. Spin-Transfer Torque (STT) or spintronic memristors), biological memristors, Micro Electromechanical Systems (MEMS). The switching mechanism is often correlated with the material type. For example, VCM and TCM switching devices are based on the movement of oxygen vacancies through the active layer of the material, while PCM devices are often constructed from a chalcogenide glass such as Ge-Sb-Te (GST).
I-B Event Based Systems
Event-based systems are systems where changes to the state are driven by the time of events occurring, rather than according to a pre-defined sample time. Such systems can efficiently process information that is sparse in time. Systems that process this kind of information would need to be simulated using finite difference equations, which may be both inefficient due to the need to simulate each and every time-step, and also inaccurate, with inaccuracies that compound over time due to the use of a non-infinitesimal (finite) time-step for simulation.
Our motivation for introducing an event-based simulation framework is to allow for the simulation of memristors in event-based settings, most notably, spiking neural networks implemented on hardware, in the context of the growing research area of neuromorphic computing (computing that aims to use principles from neuroscience to improve the processing of information) [7]. Indeed, the notion of event-based processing is inspired by the function of the human brain. Mmristors have been touted as potential neuromorphic computing elements, due to their ability to process information in-place, circumventing the Von Neumann bottleneck. This bottleneck refers to the separation of storage and processing and the associated high energy and computation time costs of transferring information from the ‘store’ (storage) to the ’mill’ (processor) and back during computation [8]. Circumventing the Von Neumann bottleneck offers the potential for massively parallel processing, akin to that possible with biological neurons.
It is clear that in order to implement event-based computation in hardware with memristors, it would be more natural to use models which are suited to event-based simulation settings.
I-C Volatility
One key aspect of information processing in the human brain is the split of memory into long-term and short-term, often labelled as long term potentiation (LTP) and short term potentation (STP). The state variable in the memristor model can capture long-term memory changes, while the possibility for autonomous state evolution captures short-term memory change to some degree. In order to enrichen the model’s ability to describe STP phenomena, we account for the possibility of short term memory outside of the state variable through the inclusion of a volatility state variable. This variable allows for the capture of a variety of physical phenomena in various memristive systems, including Joule heating in thin film metal oxide devices, non-linear drift in metal oxide devices, and state relaxation in phase change memories.
Several different phenomena in both metal oxide devices and phase change memories rely on the principle of autonomous state evolution. In metal oxide devices, the autonomous degradation of the conductive filaments results in the phenomenon of resistive drift. In phase change memories, the phenomenon of structural relaxation results in a change of the device’s resistance over time. As a result, we also extend the modelling framework by introducing a volatility state variable, as a unified way of describing the autonomous state evolution behaviour seen in a variety of memristive systems.
I-D Contributions
We introduce two major contributions to memristive modelling. Our first contribution is to look at a new simulation paradigm for event-based systems which is natively event-driven, based on the evaluation of time-to-fire for multiple parallel Poisson point processes, while also allowing for continuous-time changes - such as changes in temperature or bias - to be incorporated. We extend an existing generic memristor model, the Generalized Metastable Switch Model (GMSM) [1, 9] to the event-based setting, in order to address problems associated with the time-stepped simulation, where error is inevitable and proportional to the size of the timestep. We modify the rate calculations of the model, to base them on a more physically generic Boltzmann probability density function. We also generalise the model through the separation of the resistance and the state variable via a readout equation.
Our second modelling contribution is to generalise a variety of nonlinear, temporary and autonomous evolution phenomena observed in different types of memristive device - thin film Metal-Insulator-Metal devices and phase change memories - and propose the use of a single variable called the “volatility state variable” to singularly describe all these phenomena.
In addition, as part of our modelling approach, we propose the separation of resistance, or other output variable, calculations using a readout equation. This approach separates the underlying memristive state variable from the resistance, through the resistance readout equation, allowing for simplification of the model for the switching behaviour.
Following the development of the model, we provide a real resistive drift dataset for a titanium oxide memristive device and propose a linear conductance model to model the resistive drift in this type of device, based on a filament growth model. We then demonstrate the ability of the model to replicate the resistive drift phenomenon present under zero-bias.
@@ I-D Contributions
Notation We use parameters $E_{x}$ to denote an energy barrier height in terms of an electron, where $V_{x}=E_{x}/q$ is the corresponding equivalent potential difference. We use the notation $\mathcal{X}[i]$ to denote the $i$’th element in the ordered set $\mathcal{X}$.
@@ II Event-Based Metastable Switching Model
In this section, we introduce the proposed modelling framework. We first propose an extension to the Generalised Metastable Switch Model [1] - a general model for memristive devices - and examine a modified metastable switch model based on continuous time probabilistic point processes, rather than discrete time Bernoulli random variables, in order to make the model event based. We then incorporate the so-called “volatility state variable” to allow for richer modelling of physical phenomena such as autonomous state evolution, Joule heating, and nonlinear drift, which in turn allows for the characterisation of certain short term memory behaviours in memristive devices.
II-A Limitations of Discrete-Time Stochastic Switching Models
Many models for memristive devices have been proposed; however, only a handful of these take into account the stochasticity inherent in the devices [10, 11], with the majority making deterministic assumptions.
One general memristor model that includes stochasticity as a fundamental feature of the model, is Molter & Nugent’s “Generalized Metastable Switch Model” (GMSM) [1]. It models the memristor as a number of parallel metastable switches, each allowed to take one of two different states - the “high resistance” or “low resistance” state (see Figure 2), with stochastic transitions occurring every unit time with probabilities that are dependent on the temperature and bias.
@@ II-A Limitations of Discrete-Time Stochastic Switching Models
We note that the approximation in Equation 1 is valid only for small values of $\Delta t$. If we let the value of $\Delta t$ grow large, then errors associated with discretising the timestep begin to emerge. The larger the value of $\Delta t$, the more significant a change will have occurred in the parameters used to calculate the transition probabilities, leading to modelling inaccuracies. This is a problem for any discrete-time approximation of a continuous-time problem i.e. any finite difference method for approximating a continuous-time differential equation. Error-free modelling can only be achieved in the limit as the timestep vanishes, i.e., $\Delta t\rightarrow 0$.
In order to address the discrete-time errors associated with the GMSM, as well as being motivated by the use of memristors in event-based neuromorphic computing systems, we consider modelling the state transitions not in terms of a transition probability per timestep, but rather through a (Poisson) point process, whose rate parameter dictates the frequency of switching events. This approach switches from a timestep-based evaluation to an event-based evaluation, where we model the changes to the device’s state as and when they occur. This removes the discretisation error. It also offers potential benefits from a computational efficiency perspective, since the frequency of model updates are tied to frequency of change of the memristive state. Next, we specify the model dynamics and modelling principles.
II-B State Modelling
@@
In keeping with the terminology of the GMSM, we use the terms “switches”, or “metastable switches”, to refer to the state variables of the model (the volatility is separately dealt with). For the purposes of this work, we assume that both the number of state variables ($N$) and the multiplicity of the individual state spaces are finite. Specifically, we consider the case where $k=2$ i.e. binary switching. The extension to the continuous setting is left as future work. We use $n$ to denote the number of switches in the state that is conducive to conduction, which we call the “up state”, the “high conductance state”, or the “low resistance state”. Then, $m=N-n$ denotes the number of switches in the state that is not conducive to conduction, which we call the “down state”, the “low conductance state”, or the “high resistance state”. Switching from a lower to a higher resistance state is termed “up” switching, while switching in the opposite direction is termed “down” switching, reflecting the respective increase or decrease in the resistance. Note that due to the deterministic relationship between $m$ and $n$, there is effectively only one state variable, yet we make use of both for notational convenience.
The physical mechanism motivating the use of $N$ parallel, binary, metastable switches used as a model for memristive devices could be one of the following: the number of oxygen vacancies in the bulk of the material vs. those which are attached to metal ions, where reduction has not occurred; the number of vacancies/ions that are present in the (main) conductive filament, versus those which are not, as per the popular, single dominant conductive filament model; or the number of parallel conducting pathways running through the device, vs. those that are not formed.
We detail the inclusion of an additional state variable that models the volatility in a device, motivated by physical evidence for similar phenomena in multiple resistive switching devices, in Section II-E1. This additional variable does not affect the readout equation, but may affect the switching rates. The motivation for having a single state variable that impacts the readout equation is that this is in keeping with the spirit of most resistive switching memories, whose state is thought to be closely tied to their instantaneous resistance. In addition, it has been shown that a single state variable is sufficient to explain the behaviour of various memristive devices [5, 12].
II-C Readout Equation
@@
Note that the voltage-current relationship for a memristive device may not always be instantaneously linear. Indeed, an exponential or $\sinh$ relationship is possible depending on the nature of the device being considered [1]. Different behaviour can be considered by relating the memristive state variable, $n$, to a separate quantity such as the parameter for a Shockley diode equation, or the scaling factor for a $\sinh$ voltage-current relation.
@@
Rather than requiring that the resistance be linearly linked to the state variable, we instead use a readout equation to provide flexibility in simulation. The readout equation takes as input the memristive state variable $n$, and outputs a resistance. This allows for a consistent description of the state, and allows for the separation of measurement errors from the state evolution dynamics. In Section 10, we see an example of a thresholding, non-linear readout equation, that links the state variable to the voltage-current characteristics in a non-linear way.
II-D Switching Mechanism
@@
The GMSM models the metastable switching dynamics through a set of $N$ parallel Bernoulli random variables (one for each switch) for each timestep. As noted, our modelling approach differs in that it models the switches as independent Poisson point processes happening in parallel. The $N$ independent Bernoulli random variables are then combined into two binomial random variables - or, in the continuous approximation, a normal distribution - that measure the state change for all switches in a single timestep $\Delta T$.
@@
We instead approach the modelling by replacing the Bernoulli probabilities with rates for the independent Poisson point processes. We use the convention that the subscript up refers to parameters describing the up transition (and the same, respectively, for subscript down and down transitions).
@@
The rate parameter for the Poisson process of a single metastable switch is referred to as the base rate and denoted by either $\lambda^{\prime}_{\text{up}}$ or $\lambda^{\prime}_{\text{down}}$. In each case, its value is determined by a Boltzmann activation function (see Section II-D1 below) that is dependent on the voltage bias and the temperature, as well as static parameters. To obtain the overall rate (the cumulative rate of all up or down switching events), we simply scale the base rate by the number of switches in the origin state - $m$ for $\lambda^{\prime}_{\text{up}}$, or $n$ for $\lambda^{\prime}_{\text{down}}$, giving us the cumulative up and down switching parameters, $\lambda_{\text{up}}$ and $\lambda_{\text{down}}$. We use the convention that a positive voltage makes it more probable for the device to switch to a higher resistive state, while a negative voltage makes it more probable for the device to switch to a lower resistive state.
$\lambda^{\prime}_{\text{up}}(t)$ $=\exp{\left(-\frac{E_{a}-kV(t)q/2-E_{\text{off}}/2}{k_{B}T(t)(1+\rho(t))}\right)},$ (3)
$\lambda_{\text{up}}(t)$ $=n(t)\lambda^{\prime}_{\text{up}}(t),$ (4)
where the parameters are as follows:

### v2

@@ Abstract
In this paper, we build a general modelling framework for memristors, suitable for the simulation of event-based systems such as hardware spiking neural networks, and more generally, neuromorphic computing systems. We propose an event-based modelling approach, extending and generalising an existing general model of memristors - the Generalised Metastable Switch Model (GMSM) [1]. In doing so we eliminate errors associated discrete time approximation, as well as offering potential improvements in terms of suitability for neuromorphic memristive system simulations. We then introduce a volatility state variable to allow for the unified understanding of disparate non-linear and volatile phenomena, including state relaxation, structural disruption, Joule heating, and non-linear drift in different memristive devices. We then demonstrate a practical implementation of the modelling framework, fit to a resistive drift dataset for titanium dioxide memristors, basing our fitting on a proposed linear conductance model for resistive drift in the devices. Finally, we highlight the application of the model to neuromorphic computing, through demonstrating the contribution of the volatility state variable to switching dynamics, resulting in frequency-dependent switching (for stable memristors acting as programmable synaptic weights) and the generation of action potentials (for unstable memristors, acting as spike-generators).
@@
AI
Artificial Intelligence
ECM
Electrochemical Metallization
GMSM
Generalised Metastable Switch Model
GST
Germanium Antimony Tellurium
LTP
Long Term Potentiation
MIM
Metal-Insulator-Metal
MSE
Mean Squared Error
PCM
Phase Change Memory
RRAM
Resistive Random Access Memory
STDP
Spike Timing Dependent Plasticity
STP
Short Term Potentiation
STT
Spin Transfer Torque
TCM
Thermochemical Switching Mechanism
VCM
Valence Change Mechanism
@@ I Introduction
Figure 1: The event-based modelling framework. The framework consists of three components that are design choices: the Poisson rate parameter calculation for the state transitions, the connection between the exogenous inputs and the volatility state variable, and finally the readout equation for calculating the resistance from the state. The event-based modelling strategy is natively suited for event-driven systems and the volatility state variable can be used to explain and control key behaviours in a neuromorphic computing setting.
Memristors are emerging passive electronic devices whose resistance can be both read and modified through the application of a voltage [2, 3, 4]. Growing research interest has been in part fuelled by their possible use as fundamental computational units in the field of neuromorphic computing - a computing paradigm that aims to take computational principles from neuroscience to improve the processing of information [5].
@@
The modelling of memristive devices is an important component of the development of novel architectures and systems based on these devices. Models can be classified along several dimensions: their generality, whether they are data-driven or physics based (or both), the inclusion of stochastic or non-ideal behaviour, and suitability for simulation in terms of their computational efficiency (compactness) and simplicity of description in the relevant simulation context. The last of these dimensions can be thought of as key - since the goal of a model is to enable the prediction of behaviour so as to allow for the design of strategies to utilise the behaviour of the devices to our advantage. Many deterministic memristor models have been proposed which focus on computational efficiency and compatibility with traditional electronic design software such as SPICE. However, modelling requirements for neuromorphic computing applications differ significantly. Notably, stochastic behaviour should be taken into account, modelling should be natively compatible with event-driven simulation, and the modelling approach should be naturally aligned with the design of neuromorphic systems.
In this paper, our goal is to develop a modelling framework for memristors that is suited for their use in neuromorphic computing applications. We firstly aim to extend existing modelling approaches for memristive devices to an event-based setting to allow the models to be used for event-driven simulations, such as for spiking neural networks. Secondly, we propose a simple and unified approach to modelling device volatility, and demonstrate how such an approach can provide a useful perspective for understanding existing and potential neuromorphic behaviours (such as action potential generation and potentiation) of memristive devices, through connecting the proposed volatility state variable to the phenomena of short and long term potentiation.
The paper is structured as follows: in Section II, we begin by introducing our modelling framework, adapting an existing stochastic memristive model - the GMSM [1, 6]; we introduce an event-driven simulation approach, based on the evaluation of time-to-fire for multiple parallel Poisson point processes, while also allowing for continuous-time changes - such as changes in temperature or bias - to be incorporated. We then introduce the principle of the volatility state variable as a way to generalise and unify the description of a variety of nonlinear, temporary and autonomous evolution phenomena observed in different types of memristive device. Following the development of the model, in Section III, we demonstrate an example implementation of the modelling approach, applied to titanium dioxide memristors, using a generic Boltzmann occupation function to compute switching probabilities, modelling Joule heating using a volatility state variable, and proposing a linear readout equation motivated by a proposed physical model. We present a drift dataset collected for the titanium oxide memristors, and demonstrate the fit the model parameters to the observed data. Finally, in Section IV, we provide demonstrations of ways in which the proposed modelling framework can be used in neuromorphic simulations. We introduce the notion of stable and unstable memristors, and give illustrative examples of ways in which neuromorphic behaviours - namely, frequency-dependent switching for stable memristors and the emulation of action potentials for unstable memristors - can be brought about through an understanding of the effects of volatility on switching speed and the equilibrium point, and the subsequent appropriate selection of modelling parameters.
@@
Notation We use parameters $E_{x}$ to denote the electronic energy barrier height, where $V_{x}=E_{x}/q$ is the corresponding equivalent potential difference. We use the notation $\mathcal{X}[i]$ to denote the $i$’th element of an ordered set $\mathcal{X}$.
@@ II Event-Based Metastable Switching Model
In this section, we introduce the proposed modelling framework. In Section II-A we introduce an existing general and stochastic memristor model - the GMSM [1], detailing the introduction of a readout equation to separate state variables from summary parameters of interest such as the instantaneous device resistance. We then introduce our adaption to the event-based setting, basing the simulation on continuous time probabilistic point processes, rather than discrete time Bernoulli random variables in Section II. We introduce the “volatility state variable” in Section II-C to allow for unified description of a range of physical phenomena including autonomous state evolution, Joule heating, and nonlinear drift, which in turn will allow for the understanding of certain short term memory behaviours in memristive devices (see Section IV). In Section II-D, we specify the event-based simulation algorithm and a procedure for incorporating non-event-based inputs. We conclude the section with a discussion of the proposed model, in Section II-E.
II-A Modelling the Memristive State
II-A1 Generalised Metastable Switching Model
Many models for memristive devices have been proposed; however, only a handful of these take into account the stochasticity inherent in the devices [7, 8], with the majority making deterministic assumptions. One general memristor model that includes stochasticity as a fundamental feature of the model, is Molter & Nugent’s GMSM [1]. It models the memristor as a number of parallel metastable switches, each allowed to take one of two different states - the “high resistance” or “low resistance” state (see Figure 2), with stochastic transitions occurring every unit time with probabilities that are dependent on the temperature and bias.
@@ II-A1 Generalised Metastable Switching Model
II-A2 State Variables
@@
In keeping with the terminology of the GMSM, we use the terms “switches”, or “metastable switches”, to refer to the state variables of the model (the volatility is separately dealt with). For the purposes of this work, we assume that both the number of state variables ($N$) and the multiplicity of the individual state spaces are finite. Specifically, we consider the case where $k=2$ i.e. binary switches. Thus, the multiplicity of the state space (number of possible memristive states) is $2^{N}$. We use $n$ to denote the number of switches in the state that is conducive to conduction, which we call the “up state”, the “high conductance state”, or the “low resistance state”. Then, $m=N-n$ denotes the number of switches in the state that is not conducive to conduction, which we call the “down state”, the “low conductance state”, or the “high resistance state”. Switching from a lower to a higher resistance state is termed “up” switching, while switching in the opposite direction is termed “down” switching, reflecting the respective increase or decrease in the resistance. Note that due to the deterministic relationship between $m$ and $n$, there is effectively only one state variable, yet we make use of both for notational convenience. It has been shown that a single state variable is sufficient to explain the behaviour of various memristive devices [9, 10].
II-A3 Readout Equation
@@
Note that the voltage-current relationship for a memristive device may not always be instantaneously linear. Indeed, an exponential or $\sinh$ relationship is possible depending on the nature of the device being considered [1]. Different behaviour can be engineered by relating the memristive state variable, $n$, to a separate quantity such as the parameter for a Shockley diode equation, or the scaling factor for a $\sinh$ voltage-current relation.
@@
Rather than requiring that the resistance be linearly linked to the state variable, we instead use a readout equation to provide flexibility in simulation. The readout equation takes as input the memristive state variable $n$, and outputs a resistance. This allows for a consistent description of the state, and allows for a modular approach that separates state evolution dynamics from the resistance, which can be considered a summary statistic of a more complex state. In Section III, we see an example of a thresholding, non-linear readout equation, that links the state variable to the voltage-current characteristics in a non-linear way. The role of the readout equation in the framework is visualised in Figure 1.
II-B Event-Based Switching and the Rate Parameter Model
@@
As detailed, the GMSM models the metastable switching dynamics through a set of $N$ parallel Bernoulli random variables (one for each switch) for each timestep. The $N$ independent Bernoulli random variables are then combined into two binomial random variables - or, in the continuous approximation, a normal distribution - that measure the state change for all switches in a single timestep $\Delta T$. Motivated by the use of memristors in event-based neuromorphic computing systems, we adapt the modelling of the state transitions from a timestep-based transition probability to a (Poisson) point process, whose rate parameter dictates the frequency of switching events. This approach changes from a timestep-based evaluation to an event-based evaluation, where we model the changes to the device’s state as and when they occur. This removes disctete-time simulation errors. It also offers potential benefits from a computational efficiency perspective, since the frequency of model updates are tied to frequency of change of the memristive state.
@@
We replace the Bernoulli probabilities with Poisson rates for independent Poisson point processes. We use the convention that the subscript up refers to parameters describing the up transition (and the same, respectively, for subscript down and down transitions).
@@
The rate parameter for the Poisson process of a single metastable switch is referred to as the base rate and denoted by either $\lambda^{\prime}_{\text{up}}$ or $\lambda^{\prime}_{\text{down}}$. To obtain the overall rate (the cumulative rate of all up or down switching events), we simply scale the base rate by the number of switches in the origin state - $m$ for $\lambda^{\prime}_{\text{up}}$, or $n$ for $\lambda^{\prime}_{\text{down}}$, giving us the cumulative up and down switching parameters, $\lambda_{\text{up}}$ and $\lambda_{\text{down}}$. We use the convention that a positive voltage makes it more probable for the device to switch to a higher resistive state, while a negative voltage makes it more probable for the device to switch to a lower resistive state.
The calculation of the Poisson rate parameters for the state transitions of the metastable switches in their given state(s) will be device and material-specific. The role of the Poisson rate parameter model is shown in Figure 1. Device-to-device and cycle-to-cycle variability can be easily incorporated into the modelling by varying the parameters of the rate parameter model by device instance or programming cycle, respectively.
II-C Volatility State Variable
Different types of memristive device differ significantly in their structure, material composition, and switching/conduction mechanics, but can nevertheless display similar characteristics. We hypothesise that a single state variable can be used to describe the effect of these different phenomena for neuromorphic computing applications. Such characteristics include memory-dependence in switching and conduction behaviour; self-excitatory or self-inhibitory behaviour during switching and conduction leading to non-linear switching and/or conduction; and autonomous state evolution as a result of temporary changes within the devices, such as increased temperature or disorder in the material. Examples of phenomena that exemplify these characteristics include: